When the sun starts to cross the horizon, the surface temperature was 7 degrees at 6 am. As the day wore on, the temperature went up by 2.5 degrees per hour. By sundown, at 7 pm, what was the final daytime temperature?

Answer by stanbon(48545)

When the sun starts to cross the horizon, the surface temperature was 7 degrees at 6 am. As the day wore on, the temperature went up by 2.5 degrees per hour. By sundown, at 7 pm, what was the final daytime temperature?
---------------------------
You have two points and can generate an equation but the assumption is
that the temperature continues to rise as the day goes on; that is not true.
------------------------------
(0,7degrees) (13,9.5 degrees)
slope = 2.5
intercept = 7
EQUATION:
temp = 2.5(hrs. after 6am) + 7
temp(7pm)=2.5*11 + 7 = 34.5
===================================
Cheers,
Stan H.